### v1

@@ Abstract
The accurate determination of the true redshift distributions in tomographic bins is critical for cosmological constraints from photometric surveys. The proposed redshift self-calibration method, which utilizes the photometric galaxy clustering alone, is highly convenient and avoids the challenges from incomplete or unrepresentative spectroscopic samples in external calibration. However, the imperfection of the theoretical approximation on broad bins as well as the flaw of the previous algorithm risk the accuracy and application of the method. In this paper, we propose the improved self-calibration algorithm that incorporates novel update rules, which effectively accounts for heteroskedastic weights and noisy data with negative values. The improved algorithm greatly expands the application range of self-calibration method and accurately reconstructs the redshift distributions for various mock data. Using the luminous red galaxy (LRG) sample of the Dark Energy Spectroscopic Instrument (DESI) survey, we find that the reconstructed results are comparable to the state-of-the-art external calibration. This suggests the exciting prospect of using photometric galaxy clustering to reconstruct redshift distributions in the cosmological analysis of survey data.
keywords:
Astronomy data analysis, Redshift surveys, Clustering, Large-scale structure of the universe
[…]
xx.xx.nb, xx.xx.Ky, xx.xx.Fg
@@ 1 Introduction
[…]
Several redshift calibration techniques utilizing spectroscopic sample have been proposed in the last decades, which can be broadly separated in three categories: (i) weighting the redshift distribution of reference galaxy samples, such as direct redshift calibration (DIR; [17, 18, 19, 20, 21]) and self-organizing map (SOM; [22]) based schemes [23, 24, 25, 26, 10, 11, 27]; (ii) cross-correlating with reference spectroscopic galaxy samples [28, 29, 30, 31, 32, 19, 33, 34, 35, 21, 10, 11, 36, 37, 38, 39, 27]; (iii) Bayesian hierarchical approaches [40, 41, 42, 43, 44, 45]. Despite significant advancements and extensive application, these external calibration methods are typically restricted by a number of effects from spectroscopic samples like incompleteness and sample variance, leading to potential errors that exceed the requirements of cosmological constrains [46, 47, 48, 49, 11, 16]. Substantial advancements in redshift calibration are essential for the success of both current and forthcoming photometric surveys.
@@ 2.1 Theoretical framework
$\sigma_{ij}^{gg,P}=\sqrt{\frac{1}{(2\ell+1)\Delta\ell f_{\text{sky}}}\frac{1+\delta_{ij}}{\bar{n}_{i}\bar{n}_{j}}}\ ,$ (2)
where $\Delta\ell$ is the $\ell$ bin size, $f_{\mathrm{sky}}$ is the sky fraction, $\bar{n}_{i}$ ($\bar{n}_{j}$) is the number density in photo-$z$ bin $i$ ($j$), and $\delta_{ij}$ is the Kronecker Delta. For a given $\ell$, we can rewrite the equation (1) in matrix form,
$C_{\ell}^{gg,P}=P^{T}C_{\ell}^{gg,R}P+\delta{N_{\ell}^{gg,P}}\ .$ (3)
@@
Previous studies (e.g. [50, 52, 53]) make a crucial assumption: the photometric galaxy clustering signal can be accurately obtained from the galaxy clustering in true redshift bins by using equation (3). This assumption only holds if the galaxies scattered into a given photo-$z$ bin are representative subsamples of those in the corresponding true-$z$ bins. The simulated data in [57] satisfy this condition, as the galaxies in each photo-$z$ bin are randomly selected from the true-$z$ bins according to the preset fractions. However, galaxies from true-$z$ bins are unlikely to be a fully random subsample due to the specific characteristics of the redshift distribution profile, such as a Gaussian-like shape. The impact of this weakening becomes particularly significant in the context of broad redshift bins that were commonly used in past analyses.
@@ 2.2 Non-negative matrix factorization
$\mathcal{J}=\frac{1}{2}\sum_{\ell}\left\|C_{\ell}^{gg,P}-P^{T}C_{\ell}^{gg,R}P\right\|_{F}^{2}\ ,$ (4)
@@
$\mathcal{J}=\|X-WH\|_{F}^{2}\ ,$ (6)
which is similar to equation (4). Suppose we are minimizing an objective function over matrix $W$, the multiplicative update is easily derived from the gradient $\nabla=\partial\mathcal{J}/\partial W$. The conventional multiplicative update rule can be given by
$W\leftarrow W\circ\frac{\nabla^{-}}{\nabla^{+}}=W\circ\frac{XH^{T}}{WHH^{T}}\ ,$ (7)
where the open circle dot operator $\circ$ denotes element-wise (Hadamard) multiplication, $\nabla^{+}$ and $\nabla^{-}$ are the positive and (unsigned) negative parts of the gradient, respectively, i.e. $\nabla=\nabla^{+}-\nabla^{-}$. [53] use the iterative Lagrangian solution proposed in [68] to solve the non-negative learning problem with stochasticity constraints, i.e. $\sum_{j}W_{ij}=1$ here. The objective function is nonincreasing by using the following multiplicative update rule:
$W\leftarrow W\circ\frac{\nabla^{-}\circ A+1-B}{\nabla^{+}\circ A}\ ,$ (8)
where $\nabla^{-}=\sum_{\ell}\left[V_{\ell}H_{\ell}^{T}\right]$, $\nabla^{+}=\sum_{\ell}\left[WH_{\ell}H_{\ell}^{T}\right]$, $A_{ij}=\sum_{b}W_{ib}/\nabla_{ib}^{+}$ and $B_{ij}=\sum_{b}W_{ib}\nabla_{ib}^{-}/\nabla_{ib}^{+}$. The sum of $\ell$ components is derived from the requirement of the objective function in equation (4). Besides, the “moving term” trick [69] is applied to overcome negative entries may caused by $-B$ in the numerator, which gives
$W\leftarrow W\circ\frac{\nabla^{-}\circ A+1}{\nabla^{+}\circ A+B}\ .$ (9)
We note that additional algorithmic details like update rules for $C_{\ell}^{gg,R}$ are omitted here. The algorithm, with the update rule above, has shown promising results when handling data with high SNR, given that the equation (3) is satisfied. However, as mentioned in Section 2.1, the application of this approach to realistic scenarios demands thinner bin resolutions than the current algorithm can accommodate (see Appendix Appendix.1 for more details).
@@ 2.3 Novel update rules for NMF
$\mathcal{J}=\left\|V^{1/2}\circ\left((X+Y)-(WH+Y)\right)\right\|_{F}^{2}\ .$ (10)
where $Y$ is a shift matrix. The weight $V$ has the same dimension as $X$ and a common choice is the inverse variance matrix of each matrix element $1/\sigma^{2}$. Similar to equation (7), the update rule can be derived and given by
$W\leftarrow W\circ\frac{(V\circ(X+Y))H^{T}}{(V\circ(WH+Y))H^{T}}\ .$ (11)
@@
$(V\circ Y_{\rm min})H^{T}=\left[(V\circ X)H^{T}\right]^{-}.$ (13)
The notation $[M]^{-}$ refers to a matrix containing the absolute value of the negative elements of $M$, with the formerly positive values set to zero. Correspondingly the notation $[M]^{+}=M+[M]^{-}$. Rewriting update rule (11) with these notations, the final update rule for Nearly-NMF is given by
$W\leftarrow W\circ\frac{\left[(V\circ X)H^{T}\right]^{+}}{(V\circ(WH))H^{T}+\left[(V\circ X)H^{T}\right]^{-}}\ .$ (14)
@@
$\mathcal{J}=\frac{1}{2}\sum_{\ell}\left\|V^{1/2}\circ\left(X-WH\right)\right\|_{F}^{2}\ .$ (15)
A natural and straightforward combination idea is to consider the numerator and denominator in update rule (14) as new $\nabla^{-}$ and $\nabla^{+}$ in previous update rule (8), respectively. This substitution produces our final multiplicative update rule:
$\ W\leftarrow W\circ\frac{\nabla^{-}\circ A+1-B}{\nabla^{+}\circ A}\ ,$ (16)
@@
where $A_{ij}=\sum_{b}W_{ib}/\nabla_{ib}^{+}$, $B_{ij}=\sum_{b}W_{ib}\nabla_{ib}^{-}/\nabla_{ib}^{+}$. The same substitution operation is applied to update rule (9) in case of any negative entries in the updated $W$. We have integrated this novel update rule and objective function into the self-calibration algorithm, replacing those previously utilized in the NMF iterative process, and this modification has markedly improved the robustness and effectiveness.
@@ 3.1 DESI LRG sample
Although the majority of this LRG sample will not be observed by the DESI spectroscopy, a large number spectra within have already obtained from DESI Survey Validation (SV1; [79, 80]) which are adequate for characterizing their redshift distributions. Actually, this sample is a subset of the SV1 IR sample, and IR is one of two selections for SV1 LRG sample [75]. Therefore, [70] use 22k spectroscopic redshifts of this LRG sample from the DESI SV1 data, which covering roughly $230~{}\mathrm{deg}^{2}$ ($100~{}\mathrm{deg}^{2}$ in the North and $130~{}\mathrm{deg}^{2}$ in the South), to obtain the redshift distributions of full sample and four subsamples (Bins 1-4) shown in Figure 1. We note that the original four subsamples are obtained using slightly different redshift selections from North and South (see Table 3 of [70]). Considering the redshift distribution resolution of 0.02, and to facilitate the comparison of our reconstructed results later, we define and select these four subsamples with edges $z_{p}=$ 0.40, 0.54, 0.72, 0.86, and 1.0. Even though the selected samples may exhibit minor differences from the original, these discrepancies do not meaningfully affect the analysis or the main conclusions.
@@ 3.2 Mock data
To evaluate the performance of the enhanced algorithm featuring new updated rules, we generate various mock datasets using the reference redshift distribution derived from the DESI LRG sample. We select several specific analysis configurations. Firstly, the photometric redshifts of mocks are obtained by adding the photo-$z$ errors, which have multiple components as 1% uniform distribution, 4% Gaussian distribution with $\sigma_{z}=0.1(1+z_{s})$, and 95% Gaussian distribution with (i) $\sigma_{z}=0.03(1+z_{s})$ (hereafter Low Photo-$z$ Error) or (ii) $\sigma_{z}=0.05(1+z_{s})$ (hereafter High Photo-$z$ Error). Besides, two distinct schemes are implemented to model and manage the shot noise level with equation (2) by varying the area and surface density: (i) one scheme aligns with the DESI LRG sample above (hereafter High Noise); and (ii) the other scheme is similar to [81], designed to have a lower shot noise (hereafter Low Noise), covers a smaller area but features a higher surface density, with values of $10,000~{}\mathrm{deg}^{2}$ and $2~{}\mathrm{arcmin}^{-2}$, respectively.
@@
Applying the Limber approximation, the galaxy-galaxy angular power spectrum between two tomographic bins $i$ and $j$ is given by,
$C_{ij}^{gg}(\ell)=\int_{0}^{\chi_{\mathrm{H}}}\frac{d\chi}{\chi^{2}}K_{i}(\chi)K_{j}(\chi)P_{\delta}\left(k=\frac{\ell+1/2}{\chi},z\right)\ ,$ (17)
with $K(\chi)=H(z)b_{g}(z)n(z)$. $\chi$ and $\chi_{\mathrm{H}}$ is the comoving radial distance and the co-moving horizon distance, respectively. $H(z)=dz/d\chi$ is the expansion rate at redshift $z$, $b_{g}(z)$ is the linear galaxy bias, and $n(z)$ is the redshift distribution within tomographic bin. $P_{\delta}$ is the matter power spectrum. We assume the linear galaxy biases follow $b_{g}(z)=1.5/D(z)$, where $D(z)$ is the linear growth factor normalized by $D(z=0)=1$. We adopt a fiducial $\Lambda$CDM cosmology with parameter values from the Planck 2018 TT,TE,EE+lowE+lensing+BAO results [82].
@@
One key difference from previous works is that, we directly use the galaxy-galaxy power spectra in photometric redshift bins with redshift distributions characterized by realistic photo-$z$ errors. This approach contrasts with the aforementioned potentially risky simplifications, such as using the synthetic data directly generated by the right-hand side of equation (3), or randomly sampling a subset of simulated galaxies from true-$z$ bins. The redshift range for all mock data tests is constrained between 0.1 and 1.3, given that too low number density outside this range, making it meaningless for quantitative analysis. However, for the application to the DESI LRG sample, we utilize two broad bins on both redshift edges to encompass the complete redshift range ($1.460<z<$ 1.46) in [70].

### v3

@@ Abstract
The accurate determination of the true redshift distributions in tomographic bins is critical for cosmological constraints from photometric surveys. The proposed redshift self-calibration method, which utilizes the photometric galaxy clustering alone, is highly convenient and avoids the challenges from incomplete or unrepresentative spectroscopic samples in external calibration. However, the imperfection of the theoretical approximation on broad bins as well as the flaw of the algorithm in previous work [1] risk the accuracy and application of the method. In this paper, we propose the improved self-calibration algorithm that incorporates novel update rules, which effectively accounts for heteroskedastic weights and noisy data with negative values. The improved algorithm greatly expands the application range of self-calibration method and accurately reconstructs the redshift distributions for various mock data. Using the luminous red galaxy (LRG) sample of the Dark Energy Spectroscopic Instrument (DESI) survey, we find that the reconstructed results are comparable to the state-of-the-art external calibration. This suggests the exciting prospect of using photometric galaxy clustering to reconstruct redshift distributions in the cosmological analysis of survey data.
@@ 1 Introduction
The ongoing and upcoming photometric (imaging) galaxy surveys, such as the Hyper Suprime-Cam survey (HSC; [2]), the Kilo-Degree Survey (KIDS; [3]), the Dark Energy Survey (DES; [4]), the Dark Energy Spectroscopic Instrument (DESI; [5, 6]) Legacy Imaging Surveys, Euclid [7], the China Space Station Telescope (CSST; [8]) and the Legacy Survey of Space and Time (LSST; [9]), provide strong constraints on cosmological parameters using the clustering of galaxies and weak gravitational lensing. Nevertheless, the cosmological estimates derived from extensive photometric samples of galaxies are approaching a threshold beyond which expanding the sample size alone is unlikely to yield a marked enhancement in the precision of the constraints [10, 11, 12]. Quantifying and mitigating systematic errors is becoming crucial for reducing uncertainty in cosmological analysis [13, 14]. The major one challenge is the uncertainties of redshift distributions due to limited accuracy in the estimated photometric redshifts [15, 16, 17].
Several redshift calibration techniques utilizing spectroscopic sample have been proposed in the last decades, which can be broadly separated in three categories: (i) weighting the redshift distribution of reference galaxy samples, such as direct redshift calibration (DIR; [18, 19, 20, 21, 22]) and self-organizing map (SOM; [23]) based schemes [24, 25, 26, 27, 11, 12, 28]; (ii) cross-correlating with reference spectroscopic galaxy samples [29, 30, 31, 32, 33, 20, 34, 35, 36, 22, 11, 12, 37, 38, 39, 40, 28]; (iii) Bayesian hierarchical approaches [41, 42, 43, 44, 45, 46]. Despite significant advancements and extensive application, these external calibration methods are typically restricted by a number of effects from spectroscopic samples like incompleteness and sample variance, leading to potential errors that could surpass what is acceptable for precise cosmological constraints [47, 48, 49, 50, 12, 17]. Substantial advancements in redshift calibration are essential for the success of both current and forthcoming photometric surveys.
@@ 2.1 Theoretical framework
$\sigma_{ij}^{gg,P}=\sqrt{\frac{1}{(2\ell+1)\Delta\ell f_{\text{sky}}}\frac{1+\delta_{ij}}{\bar{n}_{i}\bar{n}_{j}}}\ ,$ (2.2)
where $\Delta\ell$ is the $\ell$ bin size, $f_{\mathrm{sky}}$ is the sky fraction, $\bar{n}_{i}$ ($\bar{n}_{j}$) is the number density in photo-$z$ bin $i$ ($j$), and $\delta_{ij}$ is the Kronecker Delta. For a given $\ell$, we can rewrite the eq. (2.1) in matrix form,
$C_{\ell}^{gg,P}=P^{T}C_{\ell}^{gg,R}P+\delta{N_{\ell}^{gg,P}}\ .$ (2.3)
@@
Previous studies (e.g. [51, 53, 54]) make a crucial assumption: the photometric galaxy clustering signal can be accurately obtained from the galaxy clustering in true redshift bins by using eq. (2.3). This assumption only holds if the galaxies scattered into a given photo-$z$ bin are representative subsamples of those in the corresponding true-$z$ bins. The simulated data in [1] satisfy this condition, as the galaxies in each photo-$z$ bin are randomly selected from the true-$z$ bins according to the preset fractions. However, galaxies from true-$z$ bins are unlikely to be a fully random subsample due to the specific characteristics of the redshift distribution profile, such as a Gaussian-like shape. The impact of this weakening becomes particularly significant in the context of broad redshift bins that were commonly used in past analyses.
@@ 2.2 Non-negative matrix factorization
$\mathcal{J}=\frac{1}{2}\sum_{\ell}\left\|C_{\ell}^{gg,P}-P^{T}C_{\ell}^{gg,R}P\right\|_{F}^{2}\ ,$ (2.4)
@@
$\mathcal{J}=\|X-WH\|_{F}^{2}\ ,$ (2.6)
which is similar to eq. (2.4). Suppose we are minimizing an objective function over matrix $W$, the multiplicative update is easily derived from the gradient $\nabla=\partial\mathcal{J}/\partial W$. The conventional multiplicative update rule can be given by
$W\leftarrow W\circ\frac{\nabla^{-}}{\nabla^{+}}=W\circ\frac{XH^{T}}{WHH^{T}}\ ,$ (2.7)
where the open circle dot operator $\circ$ denotes element-wise (Hadamard) multiplication, $\nabla^{+}$ and $\nabla^{-}$ are the positive and (unsigned) negative parts of the gradient, respectively, i.e. $\nabla=\nabla^{+}-\nabla^{-}$. Ref. [54] use the iterative Lagrangian solution proposed in [68] to solve the non-negative learning problem with stochasticity constraints, i.e. $\sum_{j}W_{ij}=1$ here. The objective function is nonincreasing by using the following multiplicative update rule:
$W\leftarrow W\circ\frac{\nabla^{-}\circ A+1-B}{\nabla^{+}\circ A}\ ,$ (2.8)
where $\nabla^{-}=\sum_{\ell}\left[V_{\ell}H_{\ell}^{T}\right]$, $\nabla^{+}=\sum_{\ell}\left[WH_{\ell}H_{\ell}^{T}\right]$, $A_{ij}=\sum_{b}W_{ib}/\nabla_{ib}^{+}$ and $B_{ij}=\sum_{b}W_{ib}\nabla_{ib}^{-}/\nabla_{ib}^{+}$. The sum of $\ell$ components is derived from the requirement of the objective function in eq. (2.4). Besides, the “moving term” trick [69] is applied to overcome negative entries may caused by $-B$ in the numerator, which gives
$W\leftarrow W\circ\frac{\nabla^{-}\circ A+1}{\nabla^{+}\circ A+B}\ .$ (2.9)
We note that additional algorithmic details like update rules for $C_{\ell}^{gg,R}$ are omitted here. The algorithm, with the update rule above, has shown promising results when handling data with high SNR, given that the eq. (2.3) is satisfied. However, as mentioned in section 2.1, the application of this approach to realistic scenarios demands thinner bin resolutions than the current algorithm can accommodate (see appendix A for more details).
@@ 2.3 Novel update rules for NMF
$\mathcal{J}=\left\|V^{1/2}\circ\left((X+Y)-(WH+Y)\right)\right\|_{F}^{2}\ ,$ (2.10)
where $Y$ is a shift matrix. The weight $V$ has the same dimension as $X$ and a common choice is the inverse variance matrix of each matrix element $1/\sigma^{2}$. Similar to eq. (2.7), the update rule can be derived and given by
$W\leftarrow W\circ\frac{(V\circ(X+Y))H^{T}}{(V\circ(WH+Y))H^{T}}\ .$ (2.11)
@@
$(V\circ Y_{\rm min})H^{T}=\left[(V\circ X)H^{T}\right]^{-}.$ (2.13)
The notation $[M]^{-}$ refers to a matrix containing the absolute value of the negative elements of $M$, with the formerly positive values set to zero. Correspondingly the notation $[M]^{+}=M+[M]^{-}$. Rewriting update rule (2.11) with these notations, the final update rule for Nearly-NMF is given by
$W\leftarrow W\circ\frac{\left[(V\circ X)H^{T}\right]^{+}}{(V\circ(WH))H^{T}+\left[(V\circ X)H^{T}\right]^{-}}\ .$ (2.14)
@@
$\mathcal{J}=\frac{1}{2}\sum_{\ell}\left\|V^{1/2}\circ\left(X-WH\right)\right\|_{F}^{2}\ .$ (2.15)
A natural and straightforward combination idea is to consider the numerator and denominator in update rule (2.14) as new $\nabla^{-}$ and $\nabla^{+}$ in previous update rule (2.8), respectively. This substitution produces our final multiplicative update rule:
$\ W\leftarrow W\circ\frac{\nabla^{-}\circ A+1-B}{\nabla^{+}\circ A}\ ,$ (2.16)
@@
where $A_{ij}=\sum_{b}W_{ib}/\nabla_{ib}^{+}$, $B_{ij}=\sum_{b}W_{ib}\nabla_{ib}^{-}/\nabla_{ib}^{+}$. The same substitution operation is applied to update rule (2.9) in case of any negative entries in the updated $W$. We have integrated this novel update rule and objective function into the self-calibration algorithm, replacing those previously utilized in the NMF iterative process, and this modification has markedly improved the robustness and effectiveness.
@@ 3.1 DESI LRG sample
Although the majority of this LRG sample will not be observed by the DESI spectroscopy, a large number spectra within have already obtained from DESI Survey Validation (SV1; [79, 80]) which are adequate for characterizing their redshift distributions. Actually, this sample is a subset of the SV1 IR sample, and IR is one of two selections for SV1 LRG sample [75]. Therefore, ref. [70] use 22k spectroscopic redshifts of this LRG sample from the DESI SV1 data, which covering roughly $230~{}\mathrm{deg}^{2}$ ($100~{}\mathrm{deg}^{2}$ in the North and $130~{}\mathrm{deg}^{2}$ in the South), to obtain the redshift distributions of full sample and four subsamples (Bins 1-4) shown in figure 1. We note that the original four subsamples are obtained using slightly different redshift selections from North and South (see Table 3 of [70]). Considering the redshift distribution resolution of 0.02, and to facilitate the comparison of our reconstructed results later, we define and select these four subsamples with edges $z_{p}=$ 0.40, 0.54, 0.72, 0.86, and 1.0. Even though the selected samples may exhibit minor differences from the original, these discrepancies do not meaningfully affect the analysis or the main conclusions.
@@ 3.2 Mock data
To evaluate the performance of the enhanced algorithm featuring new updated rules, we generate various mock datasets using the reference redshift distribution derived from the DESI LRG sample. We select several specific analysis configurations. Firstly, the photometric redshifts of mocks are obtained by adding the photo-$z$ errors, which have multiple components as 1% uniform distribution, 4% Gaussian distribution with $\sigma_{z}=0.1(1+z_{s})$, and 95% Gaussian distribution with (i) $\sigma_{z}=0.03(1+z_{s})$ (hereafter Low Photo-$z$ Error) or (ii) $\sigma_{z}=0.05(1+z_{s})$ (hereafter High Photo-$z$ Error). Besides, two distinct schemes are implemented to model and manage the shot noise level with eq. (2.2) by varying the area and surface density: (i) one scheme aligns with the DESI LRG sample above (hereafter High Noise); and (ii) the other scheme is similar to [81], designed to have a lower shot noise (hereafter Low Noise), covers a smaller area but features a higher surface density, with values of $10,000~{}\mathrm{deg}^{2}$ and $2~{}\mathrm{arcmin}^{-2}$, respectively. Note that the noise levels in the two scenarios discussed are significantly higher than what is typically expected for stage-IV cosmic shear samples (e.g. LSST) whose noise should be much smaller due to high galaxy number density.
@@
Applying the Limber approximation, the galaxy-galaxy angular power spectrum between two tomographic bins $i$ and $j$ is given by
$C_{ij}^{gg}(\ell)=\int_{0}^{\chi_{\mathrm{H}}}\frac{d\chi}{\chi^{2}}K_{i}(\chi)K_{j}(\chi)P_{\delta}\left(k=\frac{\ell+1/2}{\chi},z\right)\ ,$ (3.1)
with $K(\chi)=H(z)b_{g}(z)n(z)$. $\chi$ and $\chi_{\mathrm{H}}$ is the comoving radial distance and the co-moving horizon distance, respectively. $H(z)=dz/d\chi$ is the expansion rate at redshift $z$, $b_{g}(z)$ is the linear galaxy bias, and $n(z)$ is the redshift distribution within tomographic bin. $P_{\delta}$ is the matter power spectrum. We assume the linear galaxy biases follow $b_{g}(z)=1.5/D(z)$, where $D(z)$ is the linear growth factor normalized by $D(z=0)=1$. The idealized model of galaxy bias is scale-independent, linear and deterministic, which may not fully capture the complexities of practical scenarios. We do not take into account the possible unique bias properties of outlier galaxies, which may significantly differ from the core population, such as those originating from particular galaxy types. Neglecting this difference in galaxy bias could potentially have non-negligible implications for the self-calibration method. Nevertheless, in the mock data construction, inside a given true-$z$ bin the redshift distribution profiles for each photometric redshift bin are different, indicating that the outlier galaxies from the true-$z$ bins are not a random subsample, and at least their mean redshifts are generally different. Therefore, although there are some simplifications, the galaxy biases of outlier galaxies still exhibit some distinctions from those of the main galaxy population within each true-$z$ bin. We adopt a fiducial $\Lambda$CDM cosmology with parameter values from the Planck 2018 TT,TE,EE+lowE+lensing+BAO results [82].
@@
One key difference from previous works is that, we directly use the galaxy-galaxy power spectra in photometric redshift bins with redshift distributions characterized by realistic photo-$z$ errors. This approach contrasts with the aforementioned potentially risky simplifications, such as using the synthetic data directly generated by the right-hand side of eq. (2.3), or randomly sampling a subset of simulated galaxies from true-$z$ bins. The redshift range for all mock data tests is constrained between 0.1 and 1.3, given that too low number density outside this range, making it meaningless for quantitative analysis. However, for the application to the DESI LRG sample, we utilize two broad bins on both redshift edges to encompass the complete redshift range ($1.460<z<$ 1.46) in [70].